ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46 Azimi, Parviz, University of Sistan and Baluchestan, Iran A class of hereditarily complementably ℓ_{p} Banach spaces We construct a class of separable Banach sequence spaces. A member X of this class (i) is hereditarily complementably ℓ_{p} (1 ≤ p < ∞) dual space with the Radon-Nykodym property, (ii) if Y^{⋆} = X, there are many complemented subspaces of Y isomorphic to ℓ_{q} where \frac{1}{p} + \frac{1}{q} = 1, and (iii) X is a quasireflexive Banach space with a subsymmetric basis which as a consequence of (i) is not prime. Reference : Azimi, P., Hagler, J. : Examples of hereditarily ℓ^{1} Banach spaces failing the Schur property. Pac. J. math. 122(1986), 287-297